The income of a student production was $32,500. The price of a student ticket was $10, and the nonstudents tickets were sold at $15 each. Three thousand tickets were sold. How many tickets of each kind were sold?
Answer by mananth(16946)


Let x be number of student tickets
and y be number of non student tickets


x + y = 3000.00


10.00 x + 15.00 y = 32500.00 .............2
Eliminate y
multiply (1)by -15.00
Multiply (2) by 1.00
-15.00 x -15.00 y = -45000.00
10.00 x 15.00 y = 32500.00
Add the two equations
-5.00 x = -12500.00
/ -5.00
x = 2500.00
plug value of x in (1)
1.00 x + 1.00 y = 3000.00
2500.00 + 1.00 y = 3000.00
1.00 y = 500.00
y = 500.00
Ans x = 2500.00
y = 500.00
2500 number of student tickets
500 number of non student tickets